OpenStudy (anonymous):

how does (tan^-1(x))/x = 1 as x approaches 0?

OpenStudy (amistre64):

tan-1 is a ratio of the value of an angle; as the ratio of the angle gets smaller; we get a one on one effect that settles into something more aking to 1

OpenStudy (amistre64):

when an angle = 0 , then tangent = 0; 

OpenStudy (amistre64):

well, i should tried to drawn up an inverse tan; but same effect

OpenStudy (amistre64):

the closer to 0 we get; the closer to x/x we get

OpenStudy (amistre64):

remember, as limits go, we dont give a hoot about the value at "0" in this case; only what the value is approaching

OpenStudy (anonymous):

yes, but i need to know algebraically how to turn tan^-1 to a different ID so that when i plug in 0 for x it equals 1

OpenStudy (amistre64):

you might need more room than a single post can provide then :)

OpenStudy (amistre64):

tan-1 is an angle; a radian between -pi/2 and pi/2

OpenStudy (amistre64):

i cant think of any effective means of turning a ratio of y/x into a suitable alternative for algebraing

OpenStudy (anonymous):

like a different trig identity?

OpenStudy (amistre64):

\[tan^{-1}(x)=y\] \[x=tan(y)\] \[tan(y)/x\ ???\]

OpenStudy (amistre64):

that might be good, since x = tan(y) :) \[tan(y)\tan(y)\ ???\]

OpenStudy (amistre64):

i had a division in there: \[\lim_{x->0}\frac{tan(y)}{tan(y)}=1\] thats the best I can think of ... which may not hold water

OpenStudy (amistre64):

its bad form; tan-1(x) = y so we cant really replace it; bummer \[\frac{y}{tan(y)}\]is a better rendition

OpenStudy (amistre64):

 cant see a comparison here either